David is twice as old as his daughter. In 14 years, David's age will be three times what his daughter's age was 6 years ago. How old are they now?
Answer by TimothyLamb(4379)

d = 2k
d + 14 = 3(k - 6)
---
d + 14 = 3(k - 6)
d + 14 = 3k - 18
2k + 14 = 3k - 18
---
kid = 32
dave = 64
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php